A tourist center is open on weekends (Friday, Saturday, and Sunday). The owner-manager hopes to improve scheduling of part-time employees by determining seasonal relatives for each of these days. Data on recent traffic at the center have been tabulated and are shown in the following table:

Week 1 2 3 4 5 6

Friday 138 148 148 148 153 158

Saturday 248 248 258 258 273 273

Sunday 158 158 168 168 168 178

a. Develop seasonal relatives for the shop using the centered moving average method. (Round your intermediate calculations and final answers to 2 decimal places.)

X⎯⎯⎯ X ¯

Friday

Saturday

Sunday

b. Develop seasonal relatives for the shop using the SA method. (Round your intermediate calculations and final answers to 4 decimal places.)

SA Index

Friday

Saturday

Sunday
Answer #1

a) Seasonal relatives for the shop using the centered moving average method.

A D E Unadjusted Adjusted Seasonal Number Dt/CMA Day Week 1 Day CMA Served (Dt) relatives Seasonal Friday Saturday Sunday Tot

EXCEL FORMULAS:

Week Day Data (Dt) CMA Dt/CMA Day Unadjusted Seasonal Adjusted Seasonal relatives
1 Friday 138 Friday =AVERAGEIF($B$2:$B$19,G2,$E$2:$E$19) =H2/AVERAGE($H$2:$H$4)
Saturday 248 =AVERAGE(C2:C4) =C3/D3 Saturday =AVERAGEIF($B$2:$B$19,G3,$E$2:$E$19) =H3/AVERAGE($H$2:$H$4)
Sunday 158 =AVERAGE(C3:C5) =C4/D4 Sunday =AVERAGEIF($B$2:$B$19,G4,$E$2:$E$19) =H4/AVERAGE($H$2:$H$4)
2 Friday 148 =AVERAGE(C4:C6) =C5/D5 Total =SUM(H2:H4) =SUM(I2:I4)
Saturday 248 =AVERAGE(C5:C7) =C6/D6
Sunday 158 =AVERAGE(C6:C8) =C7/D7
3 Friday 148 =AVERAGE(C7:C9) =C8/D8
Saturday 258 =AVERAGE(C8:C10) =C9/D9
Sunday 168 =AVERAGE(C9:C11) =C10/D10
4 Friday 148 =AVERAGE(C10:C12) =C11/D11
Saturday 258 =AVERAGE(C11:C13) =C12/D12
Sunday 168 =AVERAGE(C12:C14) =C13/D13
5 Friday 153 =AVERAGE(C13:C15) =C14/D14
Saturday 273 =AVERAGE(C14:C16) =C15/D15
Sunday 168 =AVERAGE(C15:C17) =C16/D16
6 Friday 158 =AVERAGE(C16:C18) =C17/D17
Saturday 273 =AVERAGE(C17:C19) =C18/D18
Sunday 178

b) Seasonal relatives for the shop using the SA method.

A B C E G H Seasonal Seasonal Week 5 1 Day Week 1 Week 2 Week 3 Week 4 Week 6 Average Relative 2 Friday 3 Saturday 4 Sunday 1

EXCEL FORMULAS:

Friday 138 148 148 148 153 158 =AVERAGE(B2:G2) =H2/$H$5
Saturday 248 248 258 258 273 273 =AVERAGE(B3:G3) =H3/$H$5
Sunday 158 158 168 168 168 178 =AVERAGE(B4:G4) =H4/$H$5
Overall Average = =AVERAGE(H2:H4) =AVERAGE(I2:I4)